OpenStudy (anonymous):

Solve sin(x)(sin(x)-1) = 0. A. x=pi/3+or-2pin,x=3pi/4+or-2pin B. x=+of-pin C. x=+or-pin,x=pi/2+or-2pin D. x=pi/2+or-2pin,x=3pi/2+or-2pin

OpenStudy (anonymous):

Since it's already factored, what are the two possibilities? In other words,\[\text{something}=0 \vee \text{something else}=0.\] What is the "something" and "something else"?

OpenStudy (mertsj):

If the product of factors is 0 then one of the factors must be 0 so write: \[\sin x=0\] or\[\sin x-1=0\]

OpenStudy (mertsj):

Solve each of those equations.